PRACTICE 1



1. What is the percent yield if 5.50 grams of hydrogen where used in a reaction that only produced 20.4 grams NH₃?



3H² + N² □ 2 NH₃





1
Expert's answer
2022-03-16T16:43:05-0400

Solution:

The molar mass of hydrogen gas (H2) is 2.016 g/mol

Therefore,

Moles of H2 = (5.50 g H2) × (1 mol H2 / 2.016 g H2) = 2.728 mol H2


Balanced chemical equation:

3H2 + N2 → 2NH3

According to stoichiometry:

3 mol of H2 produces 2 mol of NH3

Thus, 2.728 mol of H2 produces:

(2.728 mol H2) × (2 mol NH3 / 3 mol H2) = 1.819 mol NH3


The molar mass of NH3 is 17.031 g/mol

Therefore,

Mass of NH3 = (1.819 mol NH3) × (17.031 g NH3 / 1 mol NH3) = 30.98 g NH3


Percent yield = (Actual yield / Theoretical yield) × 100%

Actual yield of NH3 is 20.4 grams

Theoretical yield of NH3 is 30.98 grams

Therefore,

Percent yield of NH3 = (20.4 g / 30.98 g) × 100% = 68.85%

Percent yield of NH3 = 68.85%


Answer: The percent yield of NH3 is 68.85%